Draw the graph of the following  all three equation 
1)X+y=10
2)2x+5y=50
3)3y-y=9
Give me the answers please 

Grade:9

1 Answers

D Vishwaksen Reddy
3 years ago
For plotting graph of a line equation, we need two parameters.
The parameters can be :
Slope and point
Two points
Two intercepts, etc 
So I choose the two points as parameters for simplicity.
For getting two points follow this algorithm:
Step 1: first take the value of X as zero and find value of y. And you got your first point as (0, your answer of y)
Step 2: now again take the equation, put value of y as zero, find X and your second point is (your answer of X, 0)
Step 3: plot your two points in the Cartesian plane
Now draw an infinite line passing through the two points.
[I'm new here so I don't know how to pin the graph so I'm giving you the points to be plotted in the graph]
1. (0,10) and (10,0)
2. (0,10) and (25,0)
3. Your question is not clear, if it is 3x-y=9
   Then points are (0,9) and (3,0).
    And if the question was 3y-x=9
    Then points are (-9,0) and (0,3)
     Or if your question is 3y-y=9
    Then draw a line parallel to x-axis passing through point 9/2 or 4.5
{ Hope my answer helps }
🙂🙂